OpenStudy (anonymous):

Help!!! Corinne solved the equation below: 5x = 95 5x − 5 = 95 − 5 x = 90 Is Corinne's solution correct? If the solution is incorrect, explain why it is incorrect and show the correct steps to solve the equation.

OpenStudy (anonymous):

5x = 95 how would you get x by itself

OpenStudy (anonymous):

you would divide 5 on both sides

OpenStudy (anonymous):

that's right. so what does x equal

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

could you help me with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

An equation is shown below: 6(2x - 11) + 15 = 21 Write the steps you will use to solve the equation and explain each step.

OpenStudy (anonymous):

ok, expand 6(2x-11)

OpenStudy (anonymous):

Expand?

OpenStudy (anonymous):

multiply 6 into the brackets

OpenStudy (anonymous):

6(2x) + 6(-11)

OpenStudy (anonymous):

12x + -66

OpenStudy (anonymous):

12x - 66 + 15 = 21

OpenStudy (anonymous):

can you solve it from here?

OpenStudy (anonymous):

i think so. give me a min

OpenStudy (anonymous):

\[-66+15=-51\]

OpenStudy (anonymous):

12x-51=21

OpenStudy (anonymous):

ok, what's the next step?

OpenStudy (anonymous):

12x=51+12?

OpenStudy (anonymous):

21, not 12

OpenStudy (anonymous):

12x=51+21 12x=72

OpenStudy (anonymous):

and then divide 12 on both sides x=6

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

Thanks you!

OpenStudy (anonymous):

thank*